# Skeleton-OOD: An End-to-End Skeleton-Based Model for Robust Out-of-Distribution Human Action Detection

### v2

@@ Abstract
keywords:
Action Recognition , Out-of-Distribution Detection , Graph Neural Networks , Energy-based Theory
@@ 1 Introduction
Human-centric visual understanding enhances the development of intelligent machines, enabling humans to accomplish greater tasks and improve their quality of life. Deep learning-based models have shown remarkable performance in supervised human action classification tasks. In skeleton-based human action datasets, this task can be solved as a typical spatio-temporal modeling problem [1, 2, 3]. Although these methods have achieved high accuracy in supervised classification problems, they still lack consideration for out-of-distribution (OOD) issues. An important assumption in these approaches is that the distribution of the training set is similar to that of the test set. However, this can not always be achieved in practice. When straightly applying these action recognition models to recognize samples outside the distribution of the train dataset, they may exhibit overconfidence by selecting a category from the known distribution [4]. Hence, it is essential for such models to be aware of the boundaries of learning for safety concerns [5], ensuring the robustness of computer vision systems. How to deal with this OOD detection problem has attracted widespread attention in fields such as autonomous driving [6] and medical image analysis [7].
@@
Furthermore, methods like [12, 13] have attempted to achieve this goal by introducing OOD related information into the training stage. However, in practice, we cannot measure the specific categories of OOD. Similarly, when the OOD category itself expands, the overconfidence problem still remains. In summary, while these methods significantly improve the accuracy of OOD recognition, achieving this goal using an end-to-end framework without prior OOD information remains challenging.
To address the limits and challenges mentioned above, we propose an end-to-end model called Skeleton-OOD to deal with the out-of-distribution detection problem for skeleton-based action recognition. Considering that the state-of-the-art feature extractors for supervised ID classification are well-developed, we focus on exploiting the extracted features to extend their adaptability for OOD detection while maintaining robust prediction performance for the ID classification. We input graphs, joints, and temporal sequence into the spatial-temporal feature extractor to obtain feature embeddings as output. Based on this, we first creatively applied a feature activation module that utilizes ASH [14] to project the original features for fitting OOD detection in the training stage. Specifically, it filters useless information and amplifies the key feature dimensions to generate a more sensitive and accurate representation of abnormal detections. However, in its original work, the activation method was only used in a post hoc manner after the trained model, and the energy score was only used for judgment. Then, regarding this as an orthogonal aggregation to the original features, we innovatively designed a feature fusion module, which integrates the original backbone outputs and the OOD-targeted activated features to enable the model to preserve a strong recognition ability for ID data while also nicely discriminating the potential OOD existences. To support the framework with end-to-end training, we suggested a meticulously designed learning loss that combines the energy function and cross-entropy information. This approach aims to improve the model’s understanding of the energy score distribution in ID data, thereby further benefiting the accurate detection of OOD samples.
@@ 2 Related works
2.1 GCNs for Skeleton based Action Recognition
@@ 2.2 Methods for out-of-distribution detection
Research on out-of-distribution detection can be mainly categorized into two ways. First, we can define a scoring function that maps each input point to a single scalar, such that in-distribution and out-of-distribution data will have different distributions. Thus, we can identify the out-of-distribution samples by calculating this score. There are many works about how to define this kind of score for pre-trained neural networks. [8] proposed the maximum predicted softmax probability (MSP). After that, [12] proved that the log of the MSP was equivalent to a special case of the free energy score. This indicates that there are some cases in ID and OOD data with similar MSP values but different energy scores. As a result, energy score has become one of the most widely used score measuring method.
@@ 3.1 GCN-based feature extractor
Figure 2: Description of the graph construction method. (a) Physically connected edges. The nodes and edges marked in each subset correspond to the nodes and edges of the two adjacent layers of nodes in the rightmost tree structure. (b) Self-loop edges (green arrow pointing to the joints themselves). We add self-loops of each joints to help the model take into account the characteristics of the joints themselves when conducting graph convolution. (c) Fully connected inter-hierarchy edges. We assume that the inter-hierarchy relationships in each subsets can help to distinguish different actions. Thus, we add fully-connected edges in each subsets as shown in yellow lines.
Let us define $G=\left\{{g_{pc},g_{sl},g_{fc}}\right\}$ as the three types of graph. The first type of graph $g_{pc}$ is physical connected edges, which are represented by blue lines between joints in Figure 2 (a). It represents the natural physical relationships in skeleton data. The second type $g_{sl}$ is the self-loop, where nodes are connected to themselves. By incorporating self-loops, nodes are enabled to consider their own features during message passing in addition to relying on neighboring nodes’ information, which helps stabilize the training process. Additionally, self-loops can be regarded as a regularization mechanism to prevent over-fitting, making them a commonly used method in graph neural network construction[31].
@@ 3.2 Feature activation shaping and fusion
There are three types of ASH strategies: ASH-P, ASH-B, and ASH-S. As shown in Figure 3, for the input feature maps $F$, ASH-P only set values less than $t$ into zero, ASH-B reset the non-zeros by binarization, and ASH-S set them with scaling. Details of these three strategies can be found in A. The difference among them lies in the normalization strategies after pruning. In the experimental part, we show the results of all these three strategies in training. After that, we retain both the output feature maps and the pruning features:
@@
Figure 3: Process of Activation Shaping strategies. They prune the elements by pruning percentage and then conduct different strategies to assign values to the remaining elements.
@@ 3.3 OOD detection using logits
For the input logits after the classifier, we measure the energy score of each sample using Eqn. 3: If the score is smaller than $\tau$, we will recognize it as ‘unseen’ for OOD and set the element value representing the probability that the sample belongs to the unknown category to 1. Otherwise, the probability of unseen will remain 0 and others (represent specific seen classes) will be the value between 0 to 1. The output of this stage will be multi-classes. The value of parameter $\tau$ is given based on the top 10% of energy scores on the training set. The assumption behind this is that the energy score of OOD after training is less than the energy score of 90% ID data. After that, we can get a vector $\hat{Y}$ that represents the probability that the sample belongs to each class.
@@
0:   Predict action label $\hat{Y}$ for inference or value of loss function $L$ in training stage.
@@
3.4 Energy based loss function
@@ 4.1 Dataset description
    NTU-RGB+D 60 [34] & 120 [35]  [^1] https://rose1.ntu.edu.sg/dataset/actionRecognition/: The NTU-RGB+D 60 dataset (NTU60), comprising 60 motion categories and 56,880 video samples, is extensively employed in diverse action recognition tasks. The NTU-RGB+D 120 dataset (NTU120) is an extension of the NTU-RGB+D 60 dataset, expanding the total number of human action categories to 120. Both datasets encompass five data modalities: 3D skeletons (body joints), masked depth maps, full-depth maps, RGB videos, and IR data. In this study, we exclusively utilize the 3D skeleton sequence data. The 3D skeleton data includes the 3D coordinates (X, Y, Z) of 25 body joints per frame for a maximum of two subjects. The actions in these datasets are categorized into three main groups: daily actions, interactive actions, and medical conditions. Five classes are randomly selected as unseen from the NTU60 dataset and ten classes from the NTU120 dataset. Note that these unseen categories encompass samples from each of the three groups.
@@
    Dataset splitting protocol: Both our training set and validation set consist of data with known labels. Only the test set contains data from unknown categories. The seen class data in the validation set and the test set are consistent. The data volume ratio of the training set and test set/validation set in the known class data is 9:1. The number of samples in the train, validation, and test dataset is shown in Table 1.
@@ 4.2.1 Experiment settings
    Implementation details: We adopt HD-GCN [23] as the backbone to extract graph-structured skeleton data as shown in Figure 1. The batch size of data is 64. The SGD optimizer is employed with a Nesterov momentum of 0.9 and a weight decay of 0.0004. The learning rate is 0.001 for NTU60 and NTU120 and 0.01 for Kinetics400. The number of learning epochs is set to 100 for NTU60 and NTU120, and 150 for Kinetics400, with a warm-up strategy [37] applied to the first five epochs for more stable learning. All our experiments are conducted on NVIDIA GeForce RTX 3090, CUDA 11.6 + PyTorch 1.12.1.
@@ 4.3.1 Performance comparison against baselines
Table 2: Comparing the results of the OOD recognition task with other methods on the mixed OOD and ID test dataset. The best results are shown in bold.
@@ Methods
As seen from Table 2, our method achieves the best performance in most metrics on the ID-OOD mixed test dataset. On the NTU60 dataset, our method achieves an average AUROC of 85.57% for OOD detection, with the accuracy of seen classification exceeding 91% in Table 3. Compared with several post-hoc methods, the average AUROC increases by 43.48%. Based on the average three evaluation metrics on the three data sets, our method also achieves the best results. To some extent, our method alleviates the problem of overconfidence in the trained model when facing OOD samples using only an end-to-end model.
For the ID sample-only evaluation in Table 3, the best accuracy in ID multi-class classification may sometimes be achieved by other methods, especially in the NTU60 dataset. However, these methods show poor performance in OOD detection tasks (see Table 2), which means that the model still recognizes most of the data as ID. Consequently, the issue of overconfidence persists. Nevertheless, based on the average performance of the three data sets, our method still obtains the highest classification accuracy in the ID sample-only test. In summary, our method maintains the accuracy of ID data classification while also ensuring effective detection capability for OOD tasks.
@@
Table 3: Comparing the results of the ID multi-class classification task with other methods on the ID sample-only test dataset. The best results are shown in bold.
ID sample-only NTU60 NTU120 Kinetics400 Average
@@
Ours-ASH-P 91.94 0.39 89.84 0.40 26.18 0.89 69.32 0.56
Ours-ASH-S 91.22 0.42 88.98 0.43 23.22 0.90 67.81 0.58
Ours-ASH-B 91.61 0.41 89.13 0.39 25.04 0.95 68.59 0.58
@@ 4.3.3 Other loss functions for OOD detection
To evaluate the effectiveness of the proposed energy-based loss function, we compare it with another method LogitNorm [41]. Experiments use our fixed framework but training with the two different loss functions. Results for both OOD detection and ID recognition tasks are shown in Table 5 and Table 6, respectively.
Table 5: Results on ID-OOD mixed dataset trained with LogitNorm loss funcion. The best results of each column are in bold.
@@ Methods
Table 6: Results on ID sample-only test dataset with ST-GCN extractor. The best results of each column are in bold.
ID sample-only NTU60 NTU120 Kinetics400 Average
@@
LogitNorm-ASH-P 90.66 0.37 89.05 0.47 26.30 0.92 68.67 0.59
LogitNorm-ASH-S 90.39 0.59 88.17 0.64 21.39 0.93 66.76 0.65
LogitNorm-ASH-B 4.44 0.79 17.63 0.79 26.19 0.91 16.09 0.83
@@ 4.3.5 Different pruning percentages
To explore the effect of pruning percentage $p$, different values are tested in our experiments. As shown in Figure 6, we select the best-performing ASH-P strategy on the NTU60 dataset for evaluation. The size of the blue points reflects the pruning ratio $p$. The red, green, and purple points respectively represent the performance of the ReAct, DICE, and DICE+ReAct methods at their optimal ratios.
@@ 4.3.6 Different backbones of skeleton feature extractor
As stated in section 1, we focus on exploiting the extracted features to extend their adaptability for OOD detection while maintaining robust prediction performance for the supervised classes. Thus, we use the state-of-the-art supervised skeleton-based recognition model HD-GCN as feature extractor. The results demonstrate that the model has achieved strong OOD recognition performance compared to many other baselines.
Additionally, we selected the ST-GCN model, which has a similar framework but underperforms compared to HD-GCN in classic action recognition tasks, to serve as a feature extractor for comparative experiments.
Table 8: Results on ID-OOD mixed dataset with ST-GCN extractor. The best results of each column are in bold.
OOD NTU60 NTU120 Kinetics400 Average
@@ Methods
Error ↓
FPR95 ↓
AUROC ↑
ST-GCN
only
35.00 65.01 81.16 56.07 68.36 53.85 51.37 95.11 51.79 47.48 76.16 62.27
ST-GCN
ASH-P
32.47 59.95 84.45 51.68 71.49 71.93 51.90 92.73 56.89 45.35 74.72 71.09
ST-GCN
ASH-S
36.83 68.67 76.63 32.96 60.93 82.90 51.82 93.09 54.85 40.54 74.23 71.46
ST-GCN
ASH-B
43.10 81.21 75.05 32.85 60.71 86.70 51.92 92.67 57.08 42.62 78.20 72.94
Table 9: Results on ID sample-only test dataset with ST-GCN extractor. The best results of each column are in bold.
ID sample-only NTU60 NTU120 Kinetics400 Average
Methods Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓
ST-GCN only 90.33 0.53 48.11 0.81 24.03 0.92 54.16 0.75
ST-GCN-ASH-P 90.97 0.43 88.99 0.72 21.93 0.89 67.30 0.68
ST-GCN-ASH-S 90.08 0.56 88.81 0.46 21.39 0.93 66.76 0.65
ST-GCN-ASH-B 90.41 0.61 87.97 0.41 21.54 0.89 66.64 0.64
Compared with the results of Table 2 and Table 8, on the OOD detection task, the performance of ST-GCN as the backbone feature extractor is not as good as that of Skeleton-OOD. On some datasets, even the best results from ST-GCN do not surpass the worst results from Skeleton-OOD. Although replacing the feature extractor with a less effective one degrades OOD recognition, the results still generally outperform those of other baselines. The results are similar in the ID task (Table 9): ST-GCN’s performance is generally lower than that of Skeleton-OOD and most baseline methods using HD-GCN as the prior skeleton feature extractor. This aligns with findings from numerous existing studies in the field of action recognition, which show that ST-GCN is inferior to HD-GCN. These indicate that ST-GCN, as a backbone, does not perform as well as Skeleton-OOD in most indicators. This further confirms that a superior feature extractor significantly aids in improving OOD recognition.
Also, comparing the results of only ST-GCN and the below activation-added methods in Table 8 and Table 9, it can be concluded that adding activation operations after the feature extractor can indeed improve the OOD detection capability of the model. This is not only limited to HD-GCN as an extractor but also works on models with other structures like ST-GCN.

### v3

@@ 1 Introduction
Human-centric visual understanding enhances the development of intelligent machines, enabling humans to accomplish greater tasks and improve their quality of life. Deep learning-based models have shown remarkable performance in supervised human action classification tasks. In skeleton-based human action datasets, this task can be solved as a typical spatio-temporal modeling problem [1, 2, 3]. Although these methods have achieved high accuracy in supervised classification problems, they still lack consideration for out-of-distribution (OOD) issues. An important assumption in these approaches is that the distribution of the training set is similar to that of the test set. However, this can not always be achieved in practice. When straightly applying these action recognition models to recognize samples outside the distribution of the training dataset, they may exhibit overconfidence by selecting a category from the known distribution [4]. Hence, it is essential for such models to be aware of the boundaries of learning for safety concerns [5], ensuring the robustness of computer vision systems. How to deal with this OOD detection problem has attracted widespread attention in fields such as autonomous driving [6] and medical image analysis [7].
@@
Furthermore, methods like [12, 13] have attempted to achieve this goal by introducing OOD-related information into the training stage. However, in practice, we cannot measure the specific categories of OOD. Similarly, when the OOD category itself expands, the overconfidence problem still remains. In summary, while these methods significantly improve the accuracy of OOD recognition, achieving this goal using an end-to-end framework without prior OOD information remains challenging.
To address the limits and challenges mentioned above, we propose an end-to-end model called Skeleton-OOD to deal with the out-of-distribution detection problem for skeleton-based action recognition. Considering that the state-of-the-art feature extractors for supervised ID classification are well-developed, we focus on exploiting the extracted features to extend their adaptability for OOD detection while maintaining robust prediction performance for ID classification. We input graphs, joints, and temporal sequences into the spatial-temporal feature extractor to obtain feature embeddings as output. Based on this, we first creatively applied a feature activation module that utilizes ASH [14] to project the original features for fitting OOD detection in the training stage. Specifically, it filters useless information and amplifies the key feature dimensions to generate a more sensitive and accurate representation of abnormal detections. However, in its original work, the activation method was only used in a post hoc manner after the trained model, and the energy score was only used for judgment. Then, regarding this as an orthogonal aggregation to the original features, we innovatively designed a feature fusion module, which integrates the original backbone outputs and the OOD-targeted activated features to enable the model to preserve a strong recognition ability for ID data while also nicely discriminating the potential OOD existences. To support the framework with end-to-end training, we suggested a meticulously designed learning loss that combines the energy function and cross-entropy information. This approach aims to improve the model’s understanding of the energy score distribution in ID data, thereby further benefiting the accurate detection of OOD samples.
@@ 2 Related works
2.1 GCNs for Skeleton-based Action Recognition
@@ 2.2 Methods for out-of-distribution detection
Research on out-of-distribution detection can be mainly categorized into two ways. First, we can define a scoring function that maps each input point to a single scalar, such that in-distribution and out-of-distribution data will have different distributions. Thus, we can identify the out-of-distribution samples by calculating this score. There are many works about how to define this kind of score for pre-trained neural networks. [8] proposed the maximum predicted softmax probability (MSP). After that, [12] proved that the log of the MSP was equivalent to a special case of the free energy score. This indicates that there are some cases in ID and OOD data with similar MSP values but different energy scores. As a result, energy score has become one of the most widely used score measuring methods.
@@ 3.1 GCN-based feature extractor
Figure 2: Description of the graph construction method. (a) Physically connected edges. The nodes and edges marked in each subset correspond to the nodes and edges of the two adjacent layers of nodes in the rightmost tree structure. (b) Self-loop edges (green arrow pointing to the joints themselves). We add self-loops of each joint to help the model take into account the characteristics of the joints themselves when conducting graph convolution. (c) Fully connected inter-hierarchy edges. We assume that the inter-hierarchy relationships in each subset can help to distinguish different actions. Thus, we add fully-connected edges in each subset as shown in yellow lines.
Let us define $G=\left\{{g_{pc},g_{sl},g_{fc}}\right\}$ as the three types of graph. The first type of graph $g_{pc}$ is physically connected edges, which are represented by blue lines between joints in Figure 2 (a). It represents the natural physical relationships in skeleton data. The second type $g_{sl}$ is the self-loop, where nodes are connected to themselves. By incorporating self-loops, nodes are enabled to consider their own features during message passing in addition to relying on neighboring nodes’ information, which helps stabilize the training process. Additionally, self-loops can be regarded as a regularization mechanism to prevent over-fitting, making them a commonly used method in graph neural network construction[31].
@@ 3.2 Feature activation shaping and fusion
There are three types of ASH strategies: ASH-P, ASH-B, and ASH-S. As shown in Figure 3, for the input feature maps $F$, ASH-P only sets values less than $t$ into zero, ASH-B resets the non-zeros by binarization, and ASH-S sets them with scaling. Details of these three strategies can be found in A. The difference among them lies in the normalization strategies after pruning. In the experimental part, we show the results of all these three strategies in training. After that, we retain both the output feature maps and the pruning features:
@@
Figure 3: Process of Activation Shaping strategies. They prune the elements by pruning percentages and then conduct different strategies to assign values to the remaining elements.
@@ 3.3 OOD detection using logits
For the input logits after the classifier, we measure the energy score of each sample using Eqn. 3: If the score is smaller than $\tau$, we will recognize it as ‘unseen’ for OOD and set the element value representing the probability that the sample belongs to the unknown category to 1. Otherwise, the probability of unseen will remain 0, and others (representing specific seen classes) will be the value between 0 to 1. The output of this stage will be multi-classes. The value of parameter $\tau$ is given based on the top 10% of energy scores on the training set. The assumption behind this is that the energy score of OOD after training is less than the energy score of 90% ID data. After that, we can get a vector $\hat{Y}$ that represents the probability that the sample belongs to each class.
@@
0:   Predict action label $\hat{Y}$ for inference or value of loss function $L$ in the training stage.
@@
3.4 Energy-based loss function
@@ 4.1 Dataset description
    NTU-RGB+D 60 [34] & 120 [35]  [^1] https://rose1.ntu.edu.sg/dataset/actionRecognition/: The NTU-RGB+D 60 dataset (NTU60), comprising 60 motion categories and 56,880 video samples, is extensively employed in diverse action recognition tasks. The NTU-RGB+D 120 dataset (NTU120) is an extension of the NTU-RGB+D 60 dataset, expanding the total number of human action categories to 120. Both datasets encompass five data modalities: 3D skeletons (body joints), masked depth maps, full-depth maps, RGB videos, and IR data. In this study, we exclusively utilize the 3D skeleton sequence data. The 3D skeleton data includes the 3D coordinates (X, Y, Z) of 25 body joints per frame for a maximum of two subjects. The actions in these datasets are categorized into three main groups: daily actions, interactive actions, and medical conditions. 5 classes are randomly selected as unseen from the NTU60 dataset and ten classes from the NTU120 dataset. Note that these unseen categories encompass samples from each of the three groups. Before splitting, we cleaned and preprocessed the raw dataset following the prior work [23].
@@
    Dataset splitting protocol: Both the training set and validation set consist of data with known labels. In this work, we create two types of test dataset: ID-only (randomly selected from known classes) and ID-OOD mix dataset, for the aim of testing the model’s ability to classify ID samples and distinguishing OOD samples (all samples of unknown classes) from mixed samples. The seen class data in the validation set and the ID-only test set are consistent. The data volume ratio of the training set and ID-only test set/validation set in the known class data is 9:1. The number of samples in the train, validation, and test dataset is shown in Table 1.
@@ 4.2.1 Experiment settings
    Implementation details: We adopt HD-GCN [23] as the backbone to extract graph-structured skeleton data as shown in Figure 1. The batch size of data is 64. The SGD optimizer is employed with a Nesterov momentum of 0.9 and a weight decay of 0.0004. Due to the differences in dataset size and data characteristics, we also used some different hyperparameters when training on different datasets: The hidden layer dimension of MLP is 400 for NTU and 576 for Kinetics. The learning rate is 0.001 for NTU60 and NTU120, and 0.01 for Kinetics400. The number of learning epochs is set to 100 for NTU60 and NTU120, and 150 for Kinetics400, with a warm-up strategy [37] applied to the first five epochs for more stable learning. The model used in this study contains a total of 1,815,060 parameters, leading to a trained model file size of approximately 9.3 MB. The training process for this model is efficient, with each epoch requiring about 7 minutes on the NTU60 dataset. The test on the NTU60 ID-OOD mix dataset (containing 9,914 samples) takes 14 seconds in total. All our experiments are conducted on NVIDIA GeForce RTX 3090, CUDA 11.6 + PyTorch 1.12.1.
@@ 4.3.1 Performance comparison against baselines
Table 2: Comparing the results of the OOD recognition task with other methods on the ID-OOD mix test dataset. The best results are shown in bold.
@@ Methods
As seen from Table 2, our method achieves the best performance in most metrics on the OOD test dataset. On the NTU60 dataset, our method achieves an average AUROC of 85.57% for OOD detection, with the accuracy of seen classification exceeding 91% in Table 3. Compared with several post-hoc methods, the average AUROC increases by 43.48%. Based on the average results on the three data sets, our method achieves the best on almost all indicators (FPR95 and AUROC). To some extent, our method alleviates the problem of overconfidence in the trained model when facing OOD samples using only an end-to-end model.
For the ID sample-only evaluation in Table 3, the best accuracy in ID multi-class classification may sometimes be achieved by other methods, especially in the NTU60 dataset. However, these methods show poor performance in OOD detection tasks (see Table 2), which means that the model still recognizes most of the data as ID. Consequently, the issue of overconfidence persists. Nevertheless, based on the average performance of the three data sets, our method still obtains the highest classification accuracy in the ID sample-only test. In summary, our method maintains the accuracy of ID data classification while also ensuring effective detection capability for OOD tasks. This phenomenon may result from ASH-P’s zero-or-not mechanism. Our analysis indicates that, with the energy-based loss function, ASH-P produces a more pronounced peak in the energy-score distribution for ID samples and a narrower range, as seen in Figure 5 (d). In contrast, the ASH-B and ASH-S methods fill in the none-zeros with the designed “averaged” values, making it harder for the model to distinguish irrelevant information during training.
@@
Table 3: Comparing the results of the ID multi-class classification task with other methods on the ID-only test dataset. The best results are shown in bold.
ID-only NTU60 NTU120 Kinetics400 Average
@@
Ours ASH-P 91.94 0.39 89.84 0.40 26.18 0.89 69.32 0.56
Ours ASH-S 91.22 0.42 88.98 0.43 23.22 0.90 67.81 0.58
Ours ASH-B 91.61 0.41 89.13 0.39 25.04 0.95 68.59 0.58
@@ 4.3.3 Other loss functions for OOD detection
To evaluate the effectiveness of the proposed energy-based loss function, we compare it with another method LogitNorm [41]. Experiments here use our proposed workflow, with the only difference being the use of two different training loss functions. Results for both OOD detection and ID recognition tasks are shown in Table 5 and Table 6, respectively.
Table 5: Results on ID-OOD mix dataset trained with LogitNorm loss function. The best results of each column are in bold.
@@ Methods
Table 6: Results on ID-only test dataset trained with LogitNorm loss function. The best results of each column are in bold.
ID-only NTU60 NTU120 Kinetics400 Average
@@
LogitNorm ASH-P 90.66 0.37 89.05 0.47 26.30 0.92 68.67 0.59
LogitNorm ASH-S 90.39 0.59 88.17 0.64 21.39 0.93 66.76 0.65
LogitNorm ASH-B 4.44 0.79 17.63 0.79 26.19 0.91 16.09 0.83
@@ 4.3.5 Different pruning percentages
To explore the effect of pruning percentage $p$, different values are tested in our experiments. As shown in Figure 6, we select the best-performing ASH-P strategy on the NTU60 dataset for evaluation. The size of the blue points reflects the pruning percentage $p$. The red, green, and purple points respectively represent the performance of the ReAct, DICE, and DICE+ReAct methods at their optimal ratios.
@@ 4.3.6 Different backbones of skeleton feature extractor
As stated in Section 1, we focus on exploiting the extracted features to extend their adaptability for OOD detection while maintaining robust prediction performance for the supervised classes. Thus, we use the state-of-the-art supervised skeleton-based recognition model HD-GCN as a feature extractor. Additionally, we also select ST-GCN, InfoGCN [42] and EfficientGCN [43], which are similar frameworks as HD-GCN in classic action recognition tasks, to serve as a feature extractor for comparative experiments. The results demonstrate that the model with HD-GCN still outperforms many other baselines.
Table 8: Results on ID-OOD mix dataset with three different extractors. The best results of each method are in bold. “+” indicates an increase in the corresponding indicator value, while “-” indicates a decrease.
OOD NTU60 NTU120 Kinetics400
@@ Methods
ST-GCN-only
35.00 65.01 81.16 56.07 68.36 53.85 51.37 95.11 51.79
InfoGCN-only
34.43 66.80 80.29 52.39 66.94 59.07 - - -
EfficientGCN-only
35.84 67.02 79.86 54.92 67.82 58.92 - - -
ST-GCN ASH-P
-2.53 -5.06 +18.08 -4.39 +3.13 +18.08 +0.53 -2.38 +5.10
ST-GCN ASH-S
+1.83 +3.66 -4.53 -23.11 -7.43 +29.05 +0.45 -2.02 +3.06
ST-GCN ASH-B
+8.10 +16.20 -6.11 -23.22 -7.65 +32.85 +0.55 -2.44 +5.29
InfoGCN ASH-P
-1.08 -0.88 +4.60 -23.22 +1.44 +28.65 - - -
InfoGCN ASH-S
-0.23 -0.06 +2.04 -9.02 +3.34 +20.75 - - -
InfoGCN ASH-B
+0.36 +1.18 -3.36 -11.52 -0.02 +24.49 - - -
EfficientGCN ASH-P
+4.44 +5.35 +1.46 -11.00 +0.56 +22.06 - - -
EfficientGCN ASH-S
-1.02 +1.90 +0.87 -15.63 -3.53 +22.28 - - -
EfficientGCN ASH-B
+7.05 +11.50 -1.12 -16.63 -3.99 +23.27 - - -
Table 9: Results on ID-only dataset with three different extractors. The best results of each method are in bold. “+” indicates an increase in the corresponding indicator value, while “-” indicates a decrease.
ID-only NTU60 NTU120 Kinetics400
Methods Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓ Top1 ↑ Overlap ↓
ST-GCN-only 90.33 0.53 48.11 0.81 24.03 0.92
InfoGCN-only 90.95 0.51 73.64 0.76 - -
EfficientGCN-only 90.88 0.55 63.21 0.79 - -
ST-GCN ASH-P +0.64 -0.10 +40.88 -0.09 -2.10 -0.03
ST-GCN ASH-S -0.25 +0.03 +40.70 -0.35 -2.64 +0.01
ST-GCN ASH-B +0.08 +0.08 +39.86 -0.40 -2.49 -0.03
InfoGCN ASH-P +0.92 -0.04 +15.30 -0.08 - -
InfoGCN ASH-S -0.01 -0.01 +15.38 -0.08 - -
InfoGCN ASH-B -0.20 +0.07 +15.60 -0.23 - -
EfficientGCN ASH-P +0.35 -0.13 +24.71 -0.32 - -
EfficientGCN ASH-S +0.14 -0.11 +25.22 -0.27 - -
EfficientGCN ASH-B -0.09 -0.02 +25.51 -0.22 - -
It should be noted that we use the Kinetics400 dataset processed by the work ST-GCN [1], while the model InfoGCN and EfficientGCN are designed for motion stream data, so the results on this dataset are missing.
Compared with the results in Table 2 and Table 8, the three feature extractor backbones perform less effectively than Skeleton-OOD on the either OOD or ID recognition task. This result indicates that HD-GCN has a stronger capability as a feature extractor than these three. This aligns with findings from numerous existing studies in the field of action recognition, which show that HD-GCN holds the state-of-the-art performance [23]. This further confirms that a superior feature extractor significantly aids in improving OOD recognition. The reason for this phenomenon may not only be related to the design of the network structure, but more importantly, the graph structures used by the methods are different.
Also, comparing the results of each backbone and the below activation-added methods in Table 8 and Table 9, it can be concluded that adding activation operations after the feature extractor can indeed improve the OOD detection capability of the model. This is not only limited to HD-GCN as an extractor but also works on models with other structures.